OpenStudy (anonymous):

The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

if a line is parallel, that means it has the same slope! So the slope of the line that contains point Q would be m= 2/3 Now i am guessing that you want to find the equation of the line? So what you want to do is substitute point Q into this equation: y - y1 = m ( x - x1) SO using Q (1, -2) , your x1 = 1 and y1 = -2

OpenStudy (anonymous):

whats the standard form?

OpenStudy (anonymous):

srry its just that these stuff are SO hard that i have no idea what to do

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):